Can someone help me with this. I had 6 problems to solve and I am having trouble solving these three. Thank you for all that you do!
1. Determine whether the following equations have a solution or not? Justify your answer.
f)s2 - 4s + 4 = 0

g)5/6x2 - 7x - 6/5 = 0

h)7a2 + 8a + 2 = 0

Answer by stanbon(75887)

1. Determine whether the following equations have a solution or not? Justify your answer.
f)s2 - 4s + 4 = 0
g)5/6x2 - 7x - 6/5 = 0
h)7a2 + 8a + 2 = 0
-----------------------------
Check the determinant of each:
f) d = 4^2 - 4*1*4 = 0
So there are two equal Real Number solutions.
--------
g) d = 7^2 - 4*(5/6)(-6/5) = 49+4 > 0
So there are two unequal Real Number solutions.
-----------
h) d = 8^2 - 4*7*2 = 64 -56 >0
So there are two unequal Real Number solutions.
-----------------------------
Cheers,
Stan H.